Cost Information During October, McCourt Associates incurred total production costs of $107,200 for copyediting manuscripts and had the following equivalent units schedule: Units completed 320 Units in EWIP × Fraction complete: 25 × 0.6 15 Equivalent units 335 Required: 1. Calculate the cost of copyediting one manuscript for October. $ per manuscript 2. Assign costs to manuscripts completed and to EWIP and then do a cost reconciliation. Manuscripts completed $ EWIP Total assigned $ 3. What if the costs assigned to units completed and EWIP total were calculated using a unit cost of $340? Calculate the total costs assigned. $ What is the discrepancy between the costs assigned and the costs to account for? $
Answer #1

Ans;

1) Cost of one manuscript = 107200/335 = 320 per manuscript

2)

Manuscripts completed (320*320)

102400

EWIP (320*15)

4800

Total Assigned

107200

3. What if the costs assigned to units completed and EWIP total were calculated using a unit cost of $340? Calculate the total costs assigned.

Manuscripts completed = 320*340 =108800

EWIP = 320*15 = 4800

Total Cost assigned = (320*340)+(320*15) = 113,600

4) What is the discrepancy between the costs assigned and the costs to account for?

113600-107200 = 6400